In a city, it is estimated that the maximum temperature in June is normally distributed with a mean of 28º and a standard deviation of 5.4°. Calculate the number of days in this month in which it is expected to reach a maximum of between 22° and 31°.


1
Expert's answer
2021-11-03T11:56:46-0400

"\mu=28 \\

\sigma=5.4 \\

P(22<X<31) = P(X<31) -P(X<22) \\

=P(Z< \frac{31-28}{5.4}) -P(Z< \frac{22-28}{5.4}) \\

= P(Z< 0.556) -P(Z< -1.111) \\

= 0.71074 -0.13326 \\

= 0.57748"

There are 30 days in the month of June.

Expected number of days in this month in which it is expected to reach a maximum of between 22° and 31° "= 30 \times 0.5775 = 17.32"

Answer: 17 days